OpenStudy (anonymous):

What is partial pressure in simple terms, or simple explanation of effect of inert gas in equilibrium of reversible reactions?

OpenStudy (abmon98):

The total pressure in any container is always equal to the sum of individual gases used. P1+P2+P3....=P total. lets take an example if we have a tank of 10 liters and lets say it contains 4.0 moles of helium and 1.10 moles of oxygen gas at room temperature (24). total pressure inside the tank, we need to know the pressure exerted by each gas and we can find PV=nRT PHe-4*8.314*295/10.0=980KPa PO2-1.10*8.314*295/10.0=270KPa Total Pressure=PHe+PO2 Total pressure=270+980=1250KPa Mixing gases combine particles and thus increase the total moles present. now comes mole fraction ratio of moles of an indivual gas to the total number of moles n1/ntotal and this is directly proportional to P1/Ptotal as volume and temp is constant.